geerky42 (geerky42):

A bag contains 1 white chip, 2 blue chips, and 3 red chips. What is the probability of randomly selecting of two chips (without replacement) that are different colors?

OpenStudy (anonymous):

You should try drawing a probability tree. Either that, or you can use the multivariate hypergeometric distribution.

geerky42 (geerky42):

So, \(\dfrac{11}{15}\)?

geerky42 (geerky42):

I think I got it, but I'm confused how it works...

OpenStudy (kropot72):

There are three mutually exclusive events: 1) Select white first and a different color next: \[P(w)=\frac{1}{6}\times 1=\frac{1}{6}\] 2) Select a blue first and a different color next: \[P(b)=\frac{2}{6}\times \frac{4}{5}=\frac{8}{30}\] 3) Select a red first and a different color next: \[P(r)=\frac{3}{6}\times \frac{3}{5}=\frac{9}{30}\] Therefore the required probability is: \[P=\frac{1}{6}+\frac{8}{30}+\frac{9}{30}=you\ can\ calculate\]